OpenStudy (anonymous):

how to solve for x 6^x=e^(x+4)

OpenStudy (anonymous):

x log 6 =( x+4) log e ?

OpenStudy (anonymous):

i got that far but am having trouble moving on from x ln(6) = x+4)

OpenStudy (anonymous):

ln6= 1.79 1.79x= x+ 4

OpenStudy (anonymous):